# Entanglement of Inhomogeneous Free Bosons and Orthogonal Polynomials

### v1

@@ I Introduction
Entanglement is a fundamentally quantum phenomenon, with no direct classical analogue. Over the past few decades, it has emerged as a central theme across a wide range of research fields, including quantum information, statistical mechanics, condensed matter physics, and high energy physics. In particular, the study of entanglement in quantum many-body systems has provided critical insights into complex emergent phenomena [1, 2], such as quantum phase transitions [3, 3, 4], topological phases of matter [5, 6], and non-equilibrium dynamics [7, 8, 9].
@@ II.1 Definition and diagonalization
$[\phi_{j},\pi_{k}]=\mathrm{i}\delta_{jk},\quad[\phi_{j},\phi_{k}]=[\pi_{j},\pi_{k}]=0.$ (5)
@@
These new operators satisfy the canonical commutation relation, $[\varphi_{k},\varpi_{k^{\prime}}]=\mathrm{i}\delta_{kk^{\prime}}$. In these coordinates, the Hamiltonian describes $N+1$ decoupled oscillators,
@@
Importantly, the eigenvalues are positive, $\lambda_{k}>0$. For a positive mass $m>0$, the eigenvalues are strictly positive and the model is gapped. For the massless case $m=0$, the lowest eigenvalues approach 00 in the thermodynamic limit $N\to\infty$, and the model becomes gapless. Since all eigenvalues are positive, the ground state of the model is the vacuum $|0\rangle$. It satisfies $a_{k}|0\rangle=0$ for all $k=0,1,\dots,N$.
@@ II.2 Entanglement entropies from correlation matrices
$\begin{split}X_{ij}&=\frac{1}{2}[\mathcal{K}^{-1/2}]_{ij}=\sum_{k=0}^{N}\frac{1}{2\lambda_{k}}U_{ik}U_{jk}\,,\\ P_{ij}&=\frac{1}{2}[\mathcal{K}^{1/2}]_{ij}=\sum_{k=0}^{N}\frac{\lambda_{k}}{2}U_{ik}U_{jk}\,,\end{split}$ (15)
@@
Second, we define the chopped correlation matrices $X_{A},P_{A}$ as $[X_{A}]_{ij}=X_{ij}$ for $i,j\in A$, and similarly for $P_{A}$. From $X_{A},P_{A}$, one defines the matrix [19]
@@
Figure 1: Left: Entanglement entropy $S_{A}$ as a function of $N$ with subsystem size $\ell=N/2.$ In the gapless case (blue curve) the entropy diverges logarithmically with $N$ as $S_{A}\sim\frac{1}{6}\log N$, whereas in the gapped case (orange curve) it is a constant in the large-$N$ limit. Right: Entanglement entropy $S_{A}$ as a function of subsystem size $\ell$ with fixed $N=250$ in the massless case. The match between the numerical data (symbols) and the prediction of Eq. (3) with $c=1$ and $s_{0}=-0.0276$ (solid line) is extremely convincing.
@@ III.1 The model
$-\Lambda_{k}u_{x}\left(\Lambda_{k}\right)=J_{x}u_{x+1}\left(\Lambda_{k}\right)-B_{x}u_{x}\left(\Lambda_{k}\right)+J_{x-1}u_{x-1}\left(\Lambda_{k}\right)\,$ (22)
@@ III.2 Inhomogeneous potential
in the large-$N$ limit. This corresponds to the scaling of the homogeneous case, multiplied by  ${a^{*}}$, which we define as the effective central charge, $c_{\rm eff}={a^{*}}$. The term effective central charge is used here in analogy with its appearance in the context of entanglement with interface defect, as discussed in [20]. However, we do not identify a conformal invariance in the models we study (apart from the homogeneous case). We note that the above results also hold when $A=\{0,1,\dots,x\}$ and $B=\{x+1,\dots,N\}$ are adjacent at a point $x$ such that $|x-x_{0}|<\mathcal{O}(\mathcal{N}_{\rm eff})$, and hence $x$ lies in the effective massless region in the thermodynamic limit.
@@ IV.2 Potential related to the Krawtchouk polynomials
where $10<p<$ 1 is a parameter. In terms of the coupling constants of the Hamiltonian (18), the Krawtchouk case corresponds to the parameters [44]
@@
Figure 3: Left: Potential $V^{K}_{x}$ in Eq. (31) corresponding to the Krawtchouk chain with $p=1/2$ and different system size $N$. The effective massless region appears to shrink for increasing values $N$. This could look contradictory to Fig. 2, but it is simply an artifact of plotting the potential as a function of $x/N$ instead of $x$. Indeed, the effective massless region scales as $N^{1/2}$, and we represent it as a function of $x/N$. Right: Entanglement entropy of the Krawtchouk chain $S_{A}^{K}$ as a function of $N$ where $A=\{0,1,\dots,x_{0}=pN\}$ for different values of $p$. The symbols are obtained by exact diagonalization and the solid lines are given by Eq. (33) with constants which are obtained by numerical fit.
@@ IV.3 Potential related to the dual Hahn polynomials
[…]
@@
Figure 5: Left: Potential $V^{H}_{x}$ in Eq. (38) corresponding to the dual Hahn chain with $\gamma=2/5$, $\delta=3/5$ and different values of $N$. Since the massless region scales as $N$, it has a constant size as a function of $x/N$, unlike the Krawtchouk case, see Fig. 3. Right: Entanglement entropy of the dual Hahn chain $S_{A}^{H}$ as a function of $N$ with $A=\{0,1,\dots,x_{0}=N\gamma/(\gamma+\delta)\}$ for different values of $\gamma,\delta$. The symbols are obtained by exact diagonalization and the solid lines are given by Eq. (40) with constants that are obtained by numerical fit.
@@ IV.4 An infinite family of effective central charges
[…]
@@
Figure 7: Entanglement entropy of the dual Hahn chain with parameters $\gamma=pN^{b}$ and $\delta=(1-p)N^{b}$ as a function of $N$ with $p=0.4$ and different values of $b$. The symbols are obtained by exact diagonalization and the solid lines are obtained from Eq. (46) with additional constants and subleading terms obtained from numerical fit.
@@ V Conclusion
There are several points of interest for future investigations. First, it would be important to apply our method to other inhomogeneous bosonic systems. For instance, one could consider more general potentials, where no analytical calculations can be performed, or other solvable models not related to orthogonal polynomials. Second, one could study the impact of the inhomogeneous potential on other entanglement measures, such as the logarithmic negativity, to further probe the effect of inhomogeneities on the entanglement structure of quantum many-body states. Finally, one should investigate further the relation between our results and the curved-space approach, both for bosonic and fermionic systems. Generalizing our results in the case of higher-dimensional systems is also promising.

### v2

@@ I Introduction
Entanglement is a fundamentally quantum phenomenon, with no direct classical analogue. Over the past few decades, it has emerged as a central theme across a wide range of research fields, including quantum information, statistical mechanics, condensed matter physics, and high energy physics. In particular, the study of entanglement in quantum many-body systems has provided critical insights into complex emergent phenomena [1, 2], such as quantum phase transitions [3, 4], topological phases of matter [5, 6], and non-equilibrium dynamics [7, 8, 9].
@@ II.1 Definition and diagonalization
$\phi_{j},\pi_{k}]=\mathrm{i}\delta_{jk},\quad[\phi_{j},\phi_{k}]=[\pi_{j},\pi_{k}]=0.$ (5)
@@
These new operators satisfy the canonical commutation relation, $\varphi_{k},\varpi_{k^{\prime}}]=\mathrm{i}\delta_{kk^{\prime}}$. In these coordinates, the Hamiltonian describes $N+1$ decoupled oscillators,
@@
Importantly, the eigenvalues are positive, $\lambda_{k}>0$. For a positive mass $m>0$, the eigenvalues are strictly positive and the model is gapped. For the massless case $m=0$, the lowest eigenvalues approach 0 in the thermodynamic limit $N\to\infty$, and the model becomes gapless. Since all eigenvalues are positive, the ground state of the model is the vacuum $|0\rangle$. It satisfies $a_{k}|0\rangle=0$ for all $k=0,1,\dots,N$.
@@ II.2 Entanglement entropies from correlation matrices
[…]
@@
Second, we define the chopped correlation matrices $X_{A},P_{A}$ as $X_{A}]_{ij}=X_{ij}$ for $i,j\in A$, and similarly for $P_{A}$. From $X_{A},P_{A}$, one defines the matrix [19]
@@
Figure 1: Left: Entanglement entropy $S_{A}$ as a function of NN with subsystem size $\ell=N/2.$ In the gapless case (blue curve) the entropy diverges logarithmically with $N$ as $S_{A}\sim\frac{1}{6}\log N$, whereas in the gapped case (orange curve) it is a constant in the large-$N$ limit. Right: Entanglement entropy $S_{A}$ as a function of subsystem size $\ell$ with fixed $N=250$ in the massless case. The match between the numerical data (symbols) and the prediction of Eq. (3) with $c=1$ and $s_{0}=-0.0276$ (solid line) is extremely convincing.
@@ III.1 The model
$\Lambda_{k}u_{x}\left(\Lambda_{k}\right)=J_{x}u_{x+1}\left(\Lambda_{k}\right)-B_{x}u_{x}\left(\Lambda_{k}\right)+J_{x-1}u_{x-1}\left(\Lambda_{k}\right)\,$ (22)
@@ III.2 Inhomogeneous potential
in the large-$N$ limit. This corresponds to the scaling of the homogeneous case, multiplied by  $a^{*}$}, which we define as the effective central charge, $c_{\rm eff}={a^{*}}$. The term effective central charge is used here in analogy with its appearance in the context of entanglement with interface defect, as discussed in [20], and does not imply an underlying conformal invariance. In particular, we do not identify a conformal invariance in the models we study, apart from the homogeneous case. We note that the above results also hold when $A=\{0,1,\dots,x\}$ and $B=\{x+1,\dots,N\}$ are adjacent at a point $x$ such that $|x-x_{0}|<\mathcal{O}(\mathcal{N}_{\rm eff})$, and hence $x$ lies in the effective massless region in the thermodynamic limit.
@@ IV.2 Potential related to the Krawtchouk polynomials
where $0<p<1$ is a parameter. In terms of the coupling constants of the Hamiltonian (18), the Krawtchouk case corresponds to the parameters [44]
@@
Figure 3: Left: Potential $V^{K}_{x}$ in Eq. (31) corresponding to the Krawtchouk chain with $p=1/2$ and different system size $N$. The effective massless region appears to shrink for increasing values NN. This could look contradictory to Fig. 2, but it is simply an artifact of plotting the potential as a function of $x/N$ instead of $x$. Indeed, the effective massless region scales as $N^{1/2}$, and we represent it as a function of $x/N$. Right: Entanglement entropy of the Krawtchouk chain $S_{A}^{K}$ as a function of NN where $A=\{0,1,\dots,x_{0}=pN\}$ for different values of $p$. The symbols are obtained by exact diagonalization and the solid lines are given by Eq. (33) with constants which are obtained by numerical fit.
@@ IV.3 Potential related to the dual Hahn polynomials
$V^{H}_{x_{0}\pm yN^{a}}=\frac{y^{2}(\gamma+\delta)^{4}}{4\gamma\delta}N^{2a-2}+\mathcal{O}(N^{\max(a-2,\ 3a-3)}),$ (39)
@@
Figure 5: Left: Potential $V^{H}_{x}$ in Eq. (38) corresponding to the dual Hahn chain with $\gamma=2/5$, $\delta=3/5$ and different values of $N$. Since the massless region scales as NN, it has a constant size as a function of $x/N$, unlike the Krawtchouk case, see Fig. 3. Right: Entanglement entropy of the dual Hahn chain $S_{A}^{H}$ as a function of NN with $A=\{0,1,\dots,x_{0}=N\gamma/(\gamma+\delta)\}$ for different values of $\gamma,\delta$. The symbols are obtained by exact diagonalization and the solid lines are given by Eq. (40) with constants that are obtained by numerical fit.
@@ IV.4 An infinite family of effective central charges
$\widetilde{V}^{H}_{x_{0}\pm yN^{a}}=\frac{y^{2}}{4p(1-p)}\times\left\{\begin{array}[]{lll}N^{2a-2+b}&+\mathcal{O}(N^{\max(a-2+b,\ 3a-3+b,\ 2a+2b-3)}),&0\leqslant b<1,\\[5.0pt] N^{2a-1}&+\mathcal{O}(N^{\max(a-1,\ 3a-2)}),&b\geqslant 1,\end{array}\right$. (44)
@@
Figure 7: Entanglement entropy of the dual Hahn chain with parameters $\gamma=pN^{b}$ and $\delta=(1-p)N^{b}$ as a function of NN with $p=0.4$ and different values of $b$. The symbols are obtained by exact diagonalization and the solid lines are obtained from Eq. (46) with additional constants and subleading terms obtained from numerical fit.
@@ V Conclusion
There are several points of interest for future investigations. First, it would be important to apply our method to other inhomogeneous bosonic systems. For instance, one could consider more general potentials, where no analytical calculations can be performed, or other solvable models not related to orthogonal polynomials. Second, one could study the impact of the inhomogeneous potential on other entanglement measures, such as the logarithmic negativity, to further probe the effect of inhomogeneities on the entanglement structure of quantum many-body states. Generalizing our results in the case of higher-dimensional systems is also promising. For inhomogeneous fermionic systems, the curved-space CFT approach is a very powerful method to extract the entanglement scaling in the thermodynamic limit [16, 46, 45]. A natural question would thus be to generalize the approach and apply it to inhomogeneous bosonic systems in order to refine the results we obtained in this paper, and in particular access the constant and subleading terms. However, this appears to be non trivial. Technically, the difficulty arises from the fact that the inhomogeneity coefficients $J_{x},V_{x}$ in the Hamiltonian (18) only pertain to the fields $\phi_{x}$, whereas the terms involving $\pi_{x}$ are homogeneous. In the continuous limit, this difference in the two fields generates consistency relations that can only be satisfied for $J_{x}=1$, i.e., the homogeneous case with an inhomogeneous potential. We leave this important point for future investigations.